# Less is more - the dispatcher/ executor principle for multi-task Reinforcement Learning

### v1

@@ 5 Discussion
This position paper introduces the idea of the dispatcher/ executor principle for structuring control architectures for multi-task reinforcement learning. It suggests a) the separation into a dispatcher module that contains general world knowledge and understands the task, b) an executor module that learns the specific interaction with a particular device and c) a regularizing communication channel between these two modules that allows for abstract and compositional communication.
We proposed a concrete implementation of such a D/E architecture for robotic manipulation. In two learning scenarios - Reinforcement learning from experience and hindsight transfer in a teacher/ student setup - we showed a massive boost in generalisation. Using the same amount of interactions with the device, the D/E architecture is able to transfer its control abilities to a whole class of related tasks.
While in this prototypical implementation many features are still engineered, next steps will concentrate on end-to-end learning architectures based on the D/E principle. The separation of the control architecture on the basis of the type of knowledge required in each part makes the dispatcher an obvious candidate to integrate large multi-modal models into the design. Their capabilities of representing general world knowledge and executing tools is what is needed to get from a task description to a tailored call of an executor. Another subject of future research is learning to find regularized representations that allow for interesting abstractions in the communication channel between dispatcher and executor.

### v2

@@ 5 Discussion
This position paper introduces the Dispatcher/Executor (D/E) principle for structuring control architectures in multi-task reinforcement learning. It advocates for: a) separating the architecture into a Dispatcher module, which holds general world knowledge and understands the task; b) an Executor module, which learns specific interactions with a particular device; and c) a regularizing communication channel that enforces abstract and compositional messaging between them.
We presented a concrete implementation of such a D/E architecture for robotic manipulation. In two distinct learning scenarios—reinforcement learning from experience and hindsight transfer—we demonstrated a substantial improvement in generalization. Using the same budget of device interactions, the D/E architecture successfully transferred its control abilities to an entire class of related tasks.
While our current implementation relies on certain engineered features, future work will focus on end-to-end learning architectures based on the D/E principle. The separation of control based on the type of knowledge required positions the Dispatcher as an ideal candidate for integrating Large Multi-modal Models (LMMs). Their capacity for general world reasoning and tool use aligns perfectly with the need to translate high-level task descriptions into tailored Executor calls. Additionally, a key subject for future research is learning regularized representations that discover optimal abstractions for the communication channel between Dispatcher and Executor.